OpenStudy (anonymous):

Multiply. Try to recognize the type of product before multiplying. (t – 1/5)^2

OpenStudy (anonymous):

Isn't it distribute property? :/

OpenStudy (anonymous):

I'm not entirely sure..

OpenStudy (anonymous):

( 2x-1)^2 = (2x-1).(2x-1)= 4x^2-4x-1

OpenStudy (unklerhaukus):

\[(t – 1/5)(t – 1/5)=t^2-\frac{2t}{5}+\frac{1}{25}\]

OpenStudy (anonymous):

thank you!